OpenStudy (raffle_snaffle):

7. Solar-heat installations successfully reduce the utility bill 60% of the time. What is the probability that at least 9 out of 10 solar-heat installations are successful and will reduce the utility bill? a. 0.0464 b. 0.9432 c. 0.0403 d. 0.8429

OpenStudy (raffle_snaffle):

@math&ing001

OpenStudy (raffle_snaffle):

@freckles

OpenStudy (raffle_snaffle):

@dan815

OpenStudy (raffle_snaffle):

@campbell_st

OpenStudy (campbell_st):

So its a binomial probability question so \[P(9 ~out~ of ~10~ Successes) = \left(\begin{matrix}10 \\ 9\end{matrix}\right)(p)^9(q)^{10-9}\] p is the probability of success q is the probability of failure hope it helps

OpenStudy (raffle_snaffle):

q = 1 - p right? I should have known this was a binomial distribution.

OpenStudy (raffle_snaffle):

I am doing the work hold on

OpenStudy (campbell_st):

that's correct p = 0.6 q = 0.4

OpenStudy (raffle_snaffle):

I get 0.09473 I don't think that is right